Circular No. 8334

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004bl IN MCG +00-31-42 Further to IAUC 8329, H. Pugh and W. Li report the LOSS discovery,

on KAIT images taken on Apr. 29.31 and 30.28 UT, of an apparent supernova (mag ∼ 16.0) located at α = 12^{h}15^{m}12^{s}.08, δ = −3^{o}26′18′′.3 (equinox 2000.0), which is 7′′.9 east and 10′′.2 north of the nucleus of MCG +00-31-42. A KAIT image taken on Apr. 12.32 showed nothing at this position (limiting mag ∼ 19.0).

COMETS C/2003 N2, O4–O8, P1, AND P2 (SOHO) Further to IAUC 8333, K. Battams reports measurements for addi-

tional Kreutz sungrazing comets found on SOHO C3 images by J. Sachs (C/2003 O4, O5, O8), X.-m. Zhou (C/2003 O6), and M. Boschat (C/2003 O7, P1, P2). C/2003 O5 was also observed in the C2 coronagraph. The reduced astrometry by R. Kisala and orbital elements by B. G. Marsden appear on the MPECs cited below. Contrary to what was stated on IAUC 8333, C/2003 N2 was only observed in the C3 coronagraph.

Comet  2003 UT  α_{2000}  δ_{2000}  MPEC
C/2003 O4 July 18.446  7^{h}35.^{m}6  +18^{o}09′  2004-H87
C/2003 O5  20.096  7 35.0  +16 46  2004-H87
C/2003 O6  24.388  7 56.9  +18 04  2004-H88
C/2003 O7  25.904  8 05.4  +18 05  2004-H88
C/2003 O8  30.138  8 21.4  +17 16  2004-H88
C/2003 P1 Aug. 2.929  8 34.3  +16 14  2004-H88
C/2003 P2  3.763  8 37.7  +16 07  2004-H88

COMET C/2004 F4 (BRADFIELD) Further visual total-magnitude and tail-length/p.a. estimates: Apr.

25.15 UT, 4.0, 8^{o}.5 in p.a. 300^{o} (J. J. Gonzalez, Leon, Spain, naked eye); 26.14, 4.6, 2.5^{o} in p.a. 310^{o} (N. Biver, Poigny-la-foret, France, 7×50 binoculars); 27.78, 5.0, 10^{o} in p.a. 305^{o} (Y. Nagai, Nagano, Japan, 7×35 binoculars); 28.79, 4.7, 6^{o} in p.a. 300^{o} (K. Yoshimoto, Yamaguchi, Japan, 7×35 binoculars); 29.02, 5.3, 4^{o} in p.a. 303^{o} (A. Baransky, Bucha, Ukraine, 10×50 binoculars); 30.08, 5.8, 3^{o}.5 in p.a. 290^{o} (K. Hornoch, Lelekovice, Czech Republic, 10×50 binoculars).

2004 April 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green